For the following problem I have to make two equations and then solve them by using determinants. This is the problem:
Fran has 108 m of fencing to enclose her house on all four sides in the shape of a rectangle. If the width is to be one-half the length, what should be the dimensions of the rectangle in meters and feet?
I think the first equation is 2L+2W=108. I'm just having trouble making the second equation.

Answer by macston(5194)

L=2W
L-2W=0
2(L+W)=108m
2L+2W=108m